Oneofus:

Circle D has a radius of 21 millimeters and central angle, ∠EDF, which measures 60°. Find the exact length of arc EF

Oneofus:

I got 7pi

Oneofus:

*7pi mm

Hero:

@Oneofus are you planning to post the image that goes with this problem?

Oneofus:

There is no imagine included @Hero

Hero:

Okay, I assumed there was but I guess we don't need one for this problem.

Hero:



Hero:



Hero:

So for this we use the formula \(S = r\theta\) where S is the arc length in question, which in this case is EF. So, you input \(r = 21\) and \(theta = \dfrac{\pi}{3}\) since in the previous problem we figured out that \(60^{\circ}\) in radians is \(\dfrac{\pi}{3}\): \(S_{EF} = 21 \times \dfrac{\pi}{3}\)

Hero:

So \(7\pi\) mm is correct.

Oneofus:

Thank you!

Hero:

You're welcome.